OpenStudy (anonymous):

9x + 2(x − 3) = 4(x − 7)

OpenStudy (anonymous):

9x + 2(x − 3) = 4(x − 7) first distribute 9x + 2x - 6 = 4x - 28 Then combine like terms 11x - 6 = 4x - 28 7x = -22 x = -22/7

OpenStudy (lilg132):

9x + 2x - 6 = 4x - 28 11x - 6 = 4x - 28 11x - 4x = -28 + 6 7x = -22 x = -3.14